Home‎ > ‎Crew 167‎ > ‎

0414-JournalistsReport

Last night Crew 167B turned the Hab into a movie theater. Crew members enjoyed spending a relaxing time together while eating buttered popcorn and watching Step Brothers.

Although today was not as beautiful as previous days with a beautiful blue sky and calm wind, it was a great day for the crew’s mission here on the red planet. At about 10:00, Martian time, Crew member LaBarre, who has been working on the energy source since Crew 167B arrived to Mars five days ago, had the opportunity of testing her Emergency Medical Service Rover (EMRS) outside the Hab. The EMSR is a robot designed and programmed to retrieve an injured astronaut out on an EVA at Mars and bring him or her back to the Hab for medical assistance. During the test LaBarre detected some operational/mechanical issues. However, she plans correcting them and a new test, pending approval, will be performed tomorrow.

Two locations were carefully selected to perform the EMRS test. The first one was the area around the Hab which offered an uneven hard surface. During the test around the Hab some operational problems were immediately detected, however, those problems were corrected shortly. On the other hand, the second location, known as Creek Bed, with a mixture of hard and soft surface, offered a more complicated terrain. This mixture of surfaces along with the variation of the terrain elevation makes it a perfect location to test the EMSR. After collecting enough information about the performance of the rover the group returned to the Hab shortly after noon, local time. The group consisted of LaBarre, Beechner, and Villarroel. All six crew members are excited while waiting for tomorrow’s results.

After enjoying spicy jambalaya rice with sausage, at about 13:45, local time, a group consisting of Rucker, Hartman, and Canham left the Hab to conduct some geology experiments. The main goal of this EVA was to study the terrain from a geological point of view and thus collect some geology data of interest. Expediters paid close attention to sedimentary rocks as well as the regolith present in the area of two different locations: Cow Patty Field and the Little Canyon. They also performed Monsell Soil Color analysis, and used a Brunton Compass to determine strike and dip. After collection of the data, the Geology EVA group returned to the Hab.

Engineers completed their daily routine without anomalies to be reported. Greenhab officer worked on taking care of the radishes, tomatoes, and green onions. A fungi looking material was observed on the green onion’s soil.

Crew member Beechner spent the afternoon working on his individual project. He has been working hard on this project during the week and good progress is noticeable. Hartman, who along with Beechner is working on a 3D virtual reality project, has showed very good progress during the week too.

Dinner was served at about 18:45, local time. Tomorrow’s plans were discussed during dinner time.

Greetings, Otsmar J. Villarroel. Commander-in-training/Astronomer/Journalist Crew 167B